A ball is moving in a circle of radius 14.9 meters has a period of 20 seconds. To the nearest tenth of a m/s, what is its speed?
Answer #1

t = 20s , r = 14.9m

Velocity V = 2.pie.r/t = 2×3.143×14.9/20 = 4.68m/s

=) V = 4.7m/s
Answer #2

To calculate the speed of the ball moving in a circle, use the formula:

v=2πrTv = \frac{2 \pi r}{T}

Where:

  • r=14.9mr = 14.9 \, \text{m} is the radius.

  • T=20secondsT = 20 \, \text{seconds} is the period.

Substitution:

v=2×3.1416×14.920v = \frac{2 \times 3.1416 \times 14.9}{20}v=93.60420v = \frac{93.604}{20}v4.7m/sv \approx 4.7 \, \text{m/s}

  •  The speed of the ball is approximately 4.7 m/s.


answered by: Monu Kumar Gupta